### v1

@@ 1.3 Irreducible Representation Subspace Symmetry Projections
The permutation and rotationally invariant encodings mentioned previously succeed in reducing the dimension of the encoding, but we note that this reduction may indeed be too drastic, which could lead to classically simulatable approaches or simply delete too much information about the input state, hindering the model performance. We therefore show how our new framework allows for linear combinations of projections that can be utilised to introduce parameterised symmetry subspace amplification. In this model, the amount of symmetry can be continuously adjusted using a hyperparameter $\alpha$. We show that by implementing an intermediate amount of permutation symmetry for point cloud data classification, it is possible to gain higher accuracy scores than using either non-invariant encodings or the fully permutation invariant encodings suggested in previous work heredge2023permutation . A visualisation of these applications is shown in Figure 1.
1.4 Linear Combinations of Unitaries Method
@@
All results in this paper are specific cases of the LCU method described in this section. Let us define the general framework of how the LCU method works in a quantum circuit as detailed in Childs2012LCU ; Berry_2015 . An insightful tutorial on the LCU method can also be found at pennylane-tutorial . The LCU procedure allows the implementation of any operator $A$ that is itself a linear combination of $N$ unitary operators $U_{j}\in SU(2^{n})$. The operator $A$ acts on a target state that is contained in an $n$ qubit target register. The target state is denoted $\lvert\psi\rangle\in(\mathbb{C}^{2})^{\otimes n}$. We can define the operator $A$ acting on $\lvert\psi\rangle$ as
@@
$P_{\text{PREP}}=$ $\frac{1}{\sqrt{\Omega}}\sum_{j=1}^{N}\sqrt{\alpha_{j}}\lvert b_{j}\rangle\langle b_{1}\rvert+\sum_{i=2}^{2^{k}}\sum_{j=1}^{2^{k}}\delta_{i,j}\lvert b_{j}\rangle\langle b_{i}\rvert,$ (7)
@@
$P_{\text{PREP}}^{\dagger}=\frac{1}{\sqrt{\Omega}}\sum_{j=1}^{N}\sqrt{\alpha_{j}}\lvert b_{1}\rangle\langle b_{j}\rvert+\sum_{i=2}^{2^{k}}\sum_{j=1}^{2^{k}}\delta_{i,j}^{*}\lvert b_{i}\rangle\langle b_{j}\rvert.$ (10)
@@
Hence, the operator $A$, which is a linear combination of unitaries and hence may itself be non-unitary, has been applied to the state $\lvert\psi\rangle$ camps2023explicit .
@@
2 Quantum Native ResNet
2.1 Variational Quantum Circuit Model Preliminaries
@@
2.2 Quantum ResNet Implementation
The key concept of ResNet he2015deep is the introduction of residual blocks. Instead of learning a mapping $H(a)$ on some data, the model instead learns some residual function $W(a)$ such that
@@ 2.2 Quantum ResNet Implementation
where $a_{0}$ is the initial data input. In order to create a quantum version of this framework, we consider the Quantum ResNet CrognalettiQResNet that can be defined as
@@
where $W_{l}(\mathbf{\theta}_{l})$ is the variational unitary for layer $l$, $\lvert\psi_{l}\rangle$ is the output of the layer, $\lvert\psi_{l-1}\rangle$ is the input of the layer, and $\lvert\psi_{0}\rangle$ corresponds to the initial input state. This initial data input state could be quantum data $\lvert\psi_{0}\rangle$ or it could be a quantum state which encodes classical data $\lvert\psi_{0}\rangle=\lvert\phi(\mathbf{x})\rangle$, where $\mathbf{x}\in\mathbb{R}^{d}$ is a $d$-dimensional classical input data vector encoded into a quantum state through a data encoding circuit $V(\mathbf{x})$ such that $\lvert\phi(\mathbf{x})\rangle=V(\mathbf{x})\lvert 0\rangle^{\otimes n}$. Implementation of Quantum ResNet of this form involves applying an operator
@@
to the state $\lvert\psi_{l-1}\rangle$. This presents a problem as the operator is not necessarily unitary in general for all $W_{l}(\mathbf{\theta}_{l})\in SU(2^{n})$. Therefore, we proceed with implementing it via the LCU method.
To maintain the most general case, and to aid in the probability scaling of the LCU procedure, we shall consider adding some additional control into the magnitude of the skipped connections parameterised for each layer by some amount $\beta_{l}\in\mathbb{R}$. Therefore the definition of a quantum residual connection we shall use is
@@
where $W_{l}(\mathbf{\theta}_{l})$ is the $l$-th layer of the variational circuit, the constants $\beta_{l}\in\mathbb{R},\beta_{l}\in[0,1]$ can be chosen freely, the input to the layer is the state $\lvert\psi_{l-1}\rangle$ and the output of the layer is the state $\lvert\psi_{l}\rangle$. The term $\Omega^{\prime}$ is a normalisation constant. The algorithm can be applied for all layers $l\in[1,L]$ simultaneously with a fixed probability of success
@@
We use the preparation and selection operators defined in Equation 25 and Equation 28, respectively, with the LCU framework as outlined in Section 1.4.
The $P_{\text{PREP}}$ operator corresponds to performing single qubit initialisation on $L$ qubits and is clearly implementable on a quantum device. Likewise $S_{\text{SELECT}}$ corresponds to a controlled $W_{l}(\mathbf{\theta}_{l})$ gate, where $W_{l}(\mathbf{\theta}_{l})$ is defined to be unitary, which is controlled by the $l$-th qubit. As $W_{l}(\mathbf{\theta}_{l})$ is unitary, the $S_{\text{SELECT}}$ operation is also implementable on a quantum device.
@@
where $\lvert 1\rangle^{(1)}(...)$ collects terms that will be discarded and can therefore be ignored. The first ancilla qubit is now measured and the result discarded unless it is found to be in the $\lvert 0\rangle^{(1)}$ state. The probability of measuring the ancilla in the $\lvert 0\rangle^{(1)}$ state will be equal to
@@
where $\Omega^{\prime}_{1}=\sqrt{\pi_{1}}$ is the required normalisation constant. Hence, for layer $l=1$ we have shown the quantum residual connection defined in Equation 22 is implemented for the original input state $\lvert\psi_{0}\rangle$.
@@
We therefore show that the most general form of ResNet where connections can be skipped at every layer in a variational quantum circuit is probabilistically implementable. An example circuit architecture that demonstrates the process in the first two layers is shown in Figure 3. This implementation requires the number of ancilla qubits to equal $L$. For a more qubit efficient but less general version see Appendix E in which the number of ancilla used scales $\mathcal{O}(\log(L))$.
2.3 Probabilistic Scaling
@@
This is dependent on the strength of residual connection $\beta_{l}$, the quantum state of the previous step $\lvert\psi_{l-1}\rangle$ and the variational circuit $W_{l}(\mathbf{\theta}_{l})$.
Note that in general $\text{Re}(\langle\psi_{l-1}\rvert W_{l}(\mathbf{\theta}_{l})\lvert\psi_{l-1}\rangle)$ lies in the range $[-1,1]$. Therefore, the probability of success lies in the range $[1-4\beta_{l}(1-\beta_{l}),1]$. The worst-case scenario therefore occurs when $\beta_{l}=\frac{1}{2}$ as in this case the probability of success is within the range $[0,1]$ and therefore there is a chance that the algorithm cannot be implemented. However, as $\beta_{l}$ is varied to be closer to 00 or $1$ this gives an increasing lower bound of $1-4\beta_{l}(1-\beta_{l})$ for the probability of success [^1] If $\beta_{l}=1$ the probability of success equals 1, but would trivially mean that the skip connection is simply not performed. Likewise, if $\beta_{l}=0$ then the success probability is equal to $1$ but corresponds to not implementing the variational circuit $W_{l}(\mathbf{\theta}_{l})$ at all.. This means that the probability of success lower bound can be adjusted to any desired value, as shown in Figure 4, by controlling the strength of the residual connection for the layer
$\lvert\psi_{l}\rangle=\frac{1}{\Omega^{\prime}}\big{(}(1-\beta_{l})\lvert\psi_{l-1}\rangle+\beta_{l}W_{l}(\mathbf{\theta}_{l})\lvert\psi_{l-1}\rangle\big{)}.$
Figure 4: Lower bound of success probability $\pi_{l}$ for layer $l$ in the LCU quantum native ResNet, as the strength of residual connection parameter $\beta_{l}$ is adjusted. The lower bound is equal to $1-4\beta_{l}(1-\beta_{l})$.
The term $\text{Re}(\langle\psi_{l-1}\rvert W_{l}(\mathbf{\theta}_{l})\lvert\psi_{l-1}\rangle)$ determines whether we will be close to the lower bound probability $1-4\beta_{l}(1-\beta_{l})$ or the upper bound of $1$. We see that if $W_{l}(\mathbf{\theta}_{l})=\mathbb{I}$ then we reach the upper bound and if $W_{l}(\mathbf{\theta}_{l})=-\mathbb{I}$ we reach the lower bound. If restrictions were placed on the form $W_{l}(\mathbf{\theta}_{l})$ can take there is potential to increase the lower bound probability even further.
@@
Therefore, the lower bound of the overall probability of success is given by $\prod_{l}^{L}\big{(}1-4\beta_{l}(1-\beta_{l})\big{)}$ which can be adjusted to be between 00 and $1$ by varying the strength of the residual layers through the variables $\beta_{l}$.
@@
If for a given architecture the probability of success for any layer is bounded between $[\pi_{\text{low}},\pi_{\text{high}}]$. Then the probability of overall success must be bounded between $[\pi_{\text{low}}^{L},\pi_{\text{high}}^{L}]$. In the case of $\pi_{\text{low}},\pi_{\text{high}}<1$, this means that the probability will decay exponentially with $L$, which is a key drawback of this method. However, if the number of layers used is chosen to scale logarithmically with the number of qubits $L=\mathcal{O}(\log(n))$ then the algorithm can still run in time $\mathcal{O}(\text{poly}(n))$ in general, where $n$ is the number of qubits for the target state register that initially contains $\lvert\psi_{0}\rangle$. Furthermore, as discussed previously, the lower bounds can be adjusted by varying the strength of the residual connections for a given architecture.
2.4 Avoiding Barren Plateaus In Complex Models
We now provide a brief sketch of a possible avenue of avoiding barren plateaus in complex variational models. We shall consider a simple example of the two-layer case $L=2$, where there is a residual connection between the input and the first layer, but no residual connection between the first and second layer. This can be implemented by setting $\beta_{1}=\frac{1}{2}$ and $\beta_{2}=1$.
@@ 2.4 Avoiding Barren Plateaus In Complex Models
$\lvert\psi_{1}\rangle=\frac{1}{2}(\lvert\psi_{0}\rangle+W_{1}(\mathbf{\theta}_{1})\lvert\psi_{0}\rangle),$ (46)
@@
hence
$\lvert\psi_{2}\rangle=\frac{1}{2}(W_{2}(\mathbf{\theta}_{2})\lvert\psi_{0}\rangle+W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})\lvert\psi_{0}\rangle).$ (48)
@@
$A(\mathbf{\theta})_{\text{RES}}=\frac{1}{2}(W_{2}(\mathbf{\theta}_{2})+W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})).$ (49)
We now consider some quantum encoded density matrix state $\rho$, where $\rho\equiv\lvert\psi_{0}\rangle\langle\psi_{0}\rvert$, which is evolved by this variational operator $A(\mathbf{\theta})_{\text{RES}}$ where the resulting loss function for the model with respect to an operator $O$ is defined by
@@
hence
@@
$\langle O\rangle_{\rho,\mathbf{\theta}}$ $=\frac{1}{4}\Big{(}\text{Tr}(W_{2}(\mathbf{\theta}_{2})\rho W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
$+\text{Tr}(W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})\rho W_{1}(\mathbf{\theta}_{1})^{\dagger}W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
$+\text{Tr}(W_{2}(\mathbf{\theta}_{2})\rho W_{1}(\mathbf{\theta}_{1})^{\dagger}W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
$+\text{Tr}(W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})\rho W_{2}(\mathbf{\theta}_{2})^{\dagger}O)\Big{)}.$ (51)
For simplicity of notation we shall assign a label to the derivative of each of the four terms.
$\negthinspace\negthinspace\negthinspace\negthinspace X^{(\mu)}_{1}\negthinspace\equiv\negthinspace\frac{\partial}{\partial\theta_{\mu}}\negthinspace\text{Tr}(W_{2}(\mathbf{\theta}_{2})\rho W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$ (52)
$\negthinspace\negthinspace\negthinspace\negthinspace X^{(\mu)}_{2}\negthinspace\equiv\negthinspace\frac{\partial}{\partial\theta_{\mu}}\negthinspace\text{Tr}(W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1}\negthinspace)\rho W_{1}(\mathbf{\theta}_{1}\negthinspace)^{\dagger}W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$ (53)
$\negthinspace\negthinspace\negthinspace\negthinspace X^{(\mu)}_{3}\negthinspace\equiv\negthinspace\frac{\partial}{\partial\theta_{\mu}}\negthinspace\text{Tr}(W_{2}(\mathbf{\theta}_{2})\rho W_{1}(\mathbf{\theta}_{1})^{\dagger}W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$ (54)
$\negthinspace\negthinspace\negthinspace\negthinspace X^{(\mu)}_{4}\negthinspace\equiv\negthinspace\frac{\partial}{\partial\theta_{\mu}}\negthinspace\text{Tr}(W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})\rho W_{2}(\mathbf{\theta}_{2})^{\dagger}O).$ (55)
where $\theta_{\mu}\in\mathbf{\theta}$ such that we can now write the gradient as

### v2

@@ I.3 Irreducible Representation Subspace Symmetry Projections
The permutation and rotationally invariant encodings mentioned previously succeed in reducing the dimension of the encoding, but we note that this reduction may indeed be too drastic, which could lead to classically simulatable approaches or simply delete too much information about the input state, hindering the model performance. We therefore show how our new framework allows for linear combinations of projections that can be utilised to introduce parameterised symmetry subspace amplification. In this setting the dimension of the quantum state can remain exponentially large, while subspaces associated with certain symmetries can have their relative weightings adjusted. We focus on the weighting of the permutation symmetric subspace relative to all other subspaces, which can be continuously adjusted using a hyperparameter $\alpha$. We show that by implementing an intermediate amount of permutation symmetry for point cloud data classification, it is possible to gain higher accuracy scores than using either non-invariant encodings or the fully permutation invariant encodings suggested in previous work Heredge et al. (2024). A visualisation of these applications is shown in Figure 1.
I.4 Linear Combinations of Unitaries Method
@@
All results in this paper are specific cases of the LCU method described in this section. Let us define the general framework of how the LCU method works in a quantum circuit as detailed in Andrew Childs (2012); Berry et al. (2015). An insightful tutorial on the LCU method can also be found at Guala et al. . The LCU procedure allows the implementation of any operator $A$ that is itself a linear combination of $N$ unitary operators $U_{j}\in SU(2^{n})$. The operator $A$ acts on a target state that is contained in an $n$ qubit target register. The target state is denoted $\lvert\psi\rangle\in(\mathbb{C}^{2})^{\otimes n}$. We can define the operator $A$ acting on $\lvert\psi\rangle$ as
@@
$P_{\text{PREP}}=$ $\frac{1}{\sqrt{\Omega}}\sum_{j=1}^{N}\sqrt{\alpha_{j}}\lvert b_{j}\rangle\langle b_{1}\rvert+\sum_{i=2}^{2^{k}}\sum_{j=1}^{2^{k}}u_{i,j}\lvert b_{j}\rangle\langle b_{i}\rvert,$ (7)
@@
$P_{\text{PREP}}^{\dagger}=\frac{1}{\sqrt{\Omega}}\sum_{j=1}^{N}\sqrt{\alpha_{j}}\lvert b_{1}\rangle\langle b_{j}\rvert+\sum_{i=2}^{2^{k}}\sum_{j=1}^{2^{k}}u_{i,j}^{*}\lvert b_{i}\rangle\langle b_{j}\rvert.$ (10)
@@
Hence, the operator $A$, which is a linear combination of unitaries and hence may itself be non-unitary, has been applied to the state $\lvert\psi\rangle$ Camps et al. (2023).
@@
II Quantum Native ResNet
II.1 Variational Quantum Circuit Model Preliminaries
@@
II.2 Quantum ResNet Implementation
The key concept of ResNet He et al. (2015) is the introduction of residual blocks. Instead of learning a mapping $H(a)$ on some data, the model instead learns some residual function $W(a)$ such that
@@ II.2 Quantum ResNet Implementation
where $a_{0}$ is the initial data input. In order to create a quantum version of this framework, we consider the quantum ResNet Crognaletti et al. (2024, ) that can be defined as
@@
where $W_{l}(\mathbf{\theta}_{l})$ is the variational unitary for layer $l$, $\lvert\psi_{l}\rangle$ is the output of the layer, $\lvert\psi_{l-1}\rangle$ is the input of the layer, and $\lvert\psi_{0}\rangle$ corresponds to the initial input state. This initial data input state could be quantum data $\lvert\psi_{0}\rangle$ or it could be a quantum state that encodes classical data $\lvert\psi_{0}\rangle=\lvert\phi(\mathbf{x})\rangle$, where $\mathbf{x}\in\mathbb{R}^{d}$ is a $d$-dimensional classical input data vector encoded into a quantum state through a data encoding circuit $V(\mathbf{x})$ such that $\lvert\phi(\mathbf{x})\rangle=V(\mathbf{x})\lvert 0\rangle^{\otimes n}$. Implementation of quantum ResNet of this form involves applying an operator
@@
to the state $\lvert\psi_{l-1}\rangle$. This presents a problem, as the operator is not necessarily unitary in general for all $W_{l}(\mathbf{\theta}_{l})\in SU(2^{n})$. Therefore, we proceed with implementing it via the LCU method.
To maintain the most general case, we shall consider adding some additional control into the magnitude of the skipped connections parameterised for each layer by some amount $\beta_{l}\in\mathbb{R}$. Therefore the definition of a quantum residual connection we shall use is
@@
where $W_{l}(\mathbf{\theta}_{l})$ is the $l$-th layer of the variational circuit, the constants $\beta_{l}\in\mathbb{R},\beta_{l}\in[0,1]$ can be freely chosen, the input to the layer is the state $\lvert\psi_{l-1}\rangle$ and the output of the layer is the state $\lvert\psi_{l}\rangle$. The term $\Omega^{\prime}$ is a normalisation constant. The algorithm can be applied for all layers $l\in[1,L]$ simultaneously with a fixed probability of success
@@
We use the preparation and selection operators defined in Equation 25 and Equation 28, respectively, with the LCU framework as outlined in Section I.4.
The $P_{\text{PREP}}$ operator corresponds to performing single qubit initialisation on $L$ qubits and is clearly implementable on a quantum device. Likewise $S_{\text{SELECT}}$ corresponds to a controlled $W_{l}(\mathbf{\theta}_{l})$ gate, where $W_{l}(\mathbf{\theta}_{l})$ is defined to be unitary, which is controlled by the $l$-th qubit. As $W_{l}(\mathbf{\theta}_{l})$ is unitary, the $S_{\text{SELECT}}$ operation can also be implemented on a quantum device.
@@
where $\lvert 1\rangle^{(1)}(...)$ collects terms that will be discarded and can therefore be ignored. The first ancilla qubit is now measured, and the result is discarded unless it is found to be in the $\lvert 0\rangle^{(1)}$ state. The probability of measuring the ancilla in the $\lvert 0\rangle^{(1)}$ state will be equal to
@@
where $\Omega^{\prime}_{1}=\sqrt{\pi_{1}}$ is the required normalisation constant. Hence, for layer $l=1$ we have shown that the quantum residual connection defined in Equation 22 is implemented for the original input state $\lvert\psi_{0}\rangle$.
@@
We therefore show that the most general form of ResNet where connections can be skipped at every layer in a variational quantum circuit is probabilistically implementable. An example circuit architecture that demonstrates the process in the first two layers is shown in Figure 3. This implementation requires the number of ancilla qubits to equal $L$. For a more qubit efficient but less general version, see Appendix G in which the number of ancilla used scales $\mathcal{O}(\log(L))$.
II.3 Probabilistic Scaling
@@
This depends on the strength of the residual connection $\beta_{l}$, the quantum state of the previous step $\lvert\psi_{l-1}\rangle$ and the variational circuit $W_{l}(\mathbf{\theta}_{l})$.
Note that in general $\text{Re}(\langle\psi_{l-1}\rvert W_{l}(\mathbf{\theta}_{l})\lvert\psi_{l-1}\rangle)$ lies in the range $[-1,1]$. Therefore, the probability of success lies in the range $[1-4\beta_{l}(1-\beta_{l}),1]$. The worst-case scenario therefore occurs when $\beta_{l}=\frac{1}{2}$ as in this case the probability of success is within the range $[0,1]$ and therefore there is a chance that the algorithm cannot be implemented. However, as $\beta_{l}$ is varied to be closer to 00 or $1$ this gives an increasing lower bound of $1-4\beta_{l}(1-\beta_{l})$ for the probability of success (if $\beta_{l}=1$ the probability of success is equal to $1$, but would trivially mean that the skip connection is simply not performed. Likewise, if $\beta_{l}=0$ then the success probability is equal to $1$ but corresponds to not implementing the variational circuit $W_{l}(\mathbf{\theta}_{l})$ at all). This means that the lower bound of the probability of success varies with the strength of the residual connection for the layer, as shown in Figure 4.
Figure 4: Lower bound of success probability $\pi_{l}$ for layer $l$ in the LCU quantum native ResNet, as the strength of residual connection parameter $\beta_{l}$ varies. The lower bound is equal to $1-4\beta_{l}(1-\beta_{l})$.
The term $\text{Re}(\langle\psi_{l-1}\rvert W_{l}(\mathbf{\theta}_{l})\lvert\psi_{l-1}\rangle)$ determines whether we will be close to the lower bound probability $1-4\beta_{l}(1-\beta_{l})$ or the upper bound of $1$. We see that if $W_{l}(\mathbf{\theta}_{l})=\mathbb{I}$ then we reach the upper bound and if $W_{l}(\mathbf{\theta}_{l})=-\mathbb{I}$ we reach the lower bound. If restrictions were placed on the form $W_{l}(\mathbf{\theta}_{l})$ can take, there is potential to increase the lower bound probability even further.
@@
Therefore, the lower bound of the overall probability of success is given by $\prod_{l}^{L}\big{(}1-4\beta_{l}(1-\beta_{l})\big{)}$ which can be adjusted to be between 00 and $1$ by varying the strength of the residual layers through the variables $\beta_{l}$. Although we highlight that this cannot be arbitrarily adjusted, as if $\beta_{l}$ reaches 00 or $1$ this would trivially correspond to no residual connection at all, and more generally the value of $\beta_{l}$ may be decided by alternative factors for a particular model or dataset.
@@
If for a given architecture the probability of success for any layer is bounded between $[\pi_{\text{low}},\pi_{\text{high}}]$. Then the probability of overall success must be bounded between $[\pi_{\text{low}}^{L},\pi_{\text{high}}^{L}]$. In the case of $\pi_{\text{low}},\pi_{\text{high}}<1$, this means that the probability will decay exponentially with $L$, which is a key drawback of this method. However, if the number of layers used is chosen to scale logarithmically with the number of qubits $L=\mathcal{O}(\log(n))$ then the algorithm can still run in time $\mathcal{O}(\text{poly}(n))$ in general, where $n$ is the number of qubits for the target state register that initially contains $\lvert\psi_{0}\rangle$. Note that when $L=1$ we have a purely unitary model. Therefore, for $L>1$ we will have access to models that are at least as expressive as a standard unitary VQC model, with the potential to be even more expressive. Hence, even setting $L=\mathcal{O}(1),L>1$ would still allow a greater variation in expressivity of the models compared to the standard unitary case. Furthermore, as discussed previously, the lower bounds can be adjusted by varying the strength of the residual connections for a given architecture.
II.4 Avoiding Barren Plateaus with Residual Connections
Several works have been undertaken to characterise the conditions under which VQC models exhibit barren plateaus Larocca et al. (2024). A model is said to contain barren plateaus if $\forall\theta_{\mu}\in\mathbf{\theta}$ the following condition holds Larocca et al. (2022)
where $\text{Var}_{\mathbf{\theta}}$ is the variance calculated over a uniform distribution of parameters $\mathbf{\theta}$. If the model exhibits exponential gradient concentration then this makes training models difficult. The structure of the variational ansatz plays an important role in determining this value, as for a wide range of cases the loss function gradient variance will decay exponentially if the dynamical Lie algebra (DLA) of the variational circuit generators is exponential in dimension Fontana et al. (2023); Ragone et al. (2024). Note that an important assumption in these characterisations is usually that the circuits are sufficiently deep to form approximate designs, meaning that the depth of the circuit is also an important consideration regarding the presence of barren plateaus Ragone et al. (2024). Previous work has shown that using an ensemble of several shallow depth models can avoid loss function concentration Friedrich and Maziero (2024). We show that the quantum ResNet provides a method of creating exponentially large ensembles of unitary VQC models, in which shallower depth components can help avoid barren plateaus, while also providing additional non-unitary contributions to the variance. However, if these shallower layers can be efficiently simulated classically and the non-unitary terms exponentially vanish, then the entire model may be at risk of being classically simulated. This finding suggests the importance of further exploring the non-unitary components of quantum ResNet models.
II.4.1 Avoiding Barren Plateaus
By introducing residual connections between layers of the ansatz it is possible to effectively mitigate barren plateaus in the model. This is similar to the classical role of ResNet in which vanishing gradients are mitigated in deep networks by allowing the gradients to skip over portions of the overall network Marion et al. (2022); He et al. (2015). For example, take the case where a variational ansatz can be separated into two separate sections $W_{1}(\mathbf{\theta}_{1})$ followed by $W_{2}(\mathbf{\theta}_{2})$. Furthermore, assume that when combined in series the overall operator $W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})$ exhibits barren plateaus. Consider a residual connection introduced between the input and the output of the first layer $W_{1}(\mathbf{\theta}_{1})$ with residual connection strength $\beta$, and no residual connection over the second layer $W_{2}(\mathbf{\theta}_{2})$. Starting with an initial state $\lvert\psi_{0}\rangle$ we see the states produced in this circuit are given by
@@ II.4.1 Avoiding Barren Plateaus
$\lvert\psi_{1}\rangle=\frac{1}{\Omega^{\prime}}((1-\beta)\lvert\psi_{0}\rangle+\beta W_{1}(\mathbf{\theta}_{1})\lvert\psi_{0}\rangle),$ (47)
@@
where $\Omega^{\prime}$ is a normalisation constant. It follows that the final state can be written as
$\lvert\psi_{2}\rangle=\frac{1}{\Omega^{\prime}}((1-\beta)W_{2}(\mathbf{\theta}_{2})\lvert\psi_{0}\rangle+\beta W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})\lvert\psi_{0}\rangle).$ (49)
@@
$A(\mathbf{\theta})_{\text{RES}}=((1-\beta)W_{2}(\mathbf{\theta}_{2})+\beta W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})),$ (50)
and normalising the resultant state by a factor $\frac{1}{\Omega^{\prime}}$. We now consider some quantum encoded density matrix state $\rho$, where $\rho\equiv\lvert\psi_{0}\rangle\langle\psi_{0}\rvert$, which is evolved by this variational operator $A(\mathbf{\theta})_{\text{RES}}$. The resulting loss function for the model with respect to a measurement operator $O$ is defined by
@@
While $A(\mathbf{\theta})_{\text{RES}}$ is not in general unitary we can decompose its corresponding loss function into unitary evolution type terms and non-unitary evolution terms by considering the following expansion
@@
$\langle O\rangle_{\rho,\mathbf{\theta}}$ $=\frac{1}{\Omega^{\prime 2}}\Big{(}(1-\beta)^{2}\text{Tr}(W_{2}(\mathbf{\theta}_{2})\rho W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
$+\beta^{2}\text{Tr}(W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})\rho W_{1}(\mathbf{\theta}_{1})^{\dagger}W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
$+\beta(1-\beta)\text{Tr}(W_{2}(\mathbf{\theta}_{2})\rho W_{1}(\mathbf{\theta}_{1})^{\dagger}W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
$+\beta(1-\beta)\text{Tr}(W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})\rho W_{2}(\mathbf{\theta}_{2})^{\dagger}O)\Big{)}.$ (52)
where we can collect the non-unitary terms together by representing the expression as